OpenStudy (saifoo.khan):

Solve.. Solve: 64^(x – 3) = 42x

OpenStudy (anonymous):

Log both sides.

OpenStudy (saifoo.khan):

Then?

OpenStudy (anonymous):

(x-3)log64=log42x

OpenStudy (anonymous):

xlog64-3log64=log42x

OpenStudy (saifoo.khan):

Hell no. i better get something to eat.

OpenStudy (anonymous):

O.o

OpenStudy (saifoo.khan):

Im too lazy, tell me the next step

OpenStudy (anonymous):

Iam sure you can take it on after that.

OpenStudy (saifoo.khan):

DO IT

OpenStudy (anonymous):

lol saif

OpenStudy (anonymous):

Hmm..

OpenStudy (anonymous):

^ lawlz

OpenStudy (anonymous):

Nice. Now finish the problem.

OpenStudy (saifoo.khan):

Badman! wth

OpenStudy (anonymous):

What?

OpenStudy (saifoo.khan):

wth = what the helll. Do it

OpenStudy (anonymous):

Umm..Naw, thats alrtite :D

OpenStudy (anonymous):

Oh, and plz watch language, Thanks!

OpenStudy (agreene):

\[x\approx 9.08262\times 10^{-8}\] \[x\approx 4.24643\]

OpenStudy (saifoo.khan):

After this, xlog64-3log64=log42x ?

OpenStudy (anonymous):

How did you do it?

OpenStudy (anonymous):

@Seafood

OpenStudy (saifoo.khan):

do it man

OpenStudy (anonymous):

I want to know how your setup looked like.

OpenStudy (saifoo.khan):

@agreene will do it.. dont worry.

OpenStudy (agreene):

64^(x – 3) = 42x use mobius transform: x=1/21 2^(-19+6 x) use mobius transform: 64^x=11010048 x log both sides xlog64=log1101048x solve using log product function W(z)

OpenStudy (agreene):

I used CAS, but thats basically how you approach it if you didnt want a huge headache.

OpenStudy (saifoo.khan):

That's something which got over my head. hehe How can we do after this? just tell me 1 step xlog64-3log64=log42x

OpenStudy (saifoo.khan):

@agreene

OpenStudy (agreene):

I have no idea--that's not the way I would try and solve it. partially because im horrible with logs. I suppose you would want to isolate the variables so the next step would probably be: xlog64-log42x=3log64

OpenStudy (saifoo.khan):

Alright thanks.

OpenStudy (saifoo.khan):

@GOODMAN ?

OpenStudy (agreene):

http://en.wikipedia.org/wiki/Lambert_W_function might be of use.

OpenStudy (agreene):

specifically, examples/applications #4

OpenStudy (saifoo.khan):

WOAH!